# From homeostasis to resource sharing: Biologically and economically compatible multi-objective multi-agent AI safety benchmarks

### v1

@@ 3.1 Baseline agents
As baselines, we provide two agent types - “random” and “handwritten rules”, to establish approximate bounds of lowest and highest expected scores per each benchmark.
@@ 3.1.1 Random Agent (no learning)
This agent type represents the approximate lowest scores an incompetent agent is expected to get in the benchmarks (unless it actively attempts to harm its scores further, a scenario which we do not consider here). This agent does not have a capability of learning.
@@ 3.1.2 Handwritten Rules Agent (no learning)
This agent type represents the approximate best scores an agent is expected to achieve in the benchmarks. This agent is based on simulation of hand-written heuristic action rules for a particular benchmark setup. This agent does not have an actual capability of learning.
@@ 3.2 Experimental results
Since the above mentioned baseline agents do not learn, the benchmarks did not contain a training phase. We measured the performance only during the testing phase. The environment map size was 7 tiles horizontally and vertically. Of these 7x7 maps the outer ring consisted of walls, which means the inner map for actual activities had a size of 5x5 tiles.
The random agent and handwritten rules agent simulations were tested for 1000 episodes, each episode consisted of 400 steps.
Suffice to say, if throughout a statistically significant number of runs a RL agent would fail even once, there might be considerable risks in long-term deployment of said agent in the real world. In our preliminary experiments with DQN learning we saw for example cases where the agent performed notably well in 95
In the following table you can find the average score per test step for both agent types.
@@
Benchmark Random Handwritten Rules
Food Unbounded 0.76 19.9
Danger Tiles -0.9 19.85
Predators -3.1 18.95
Homeostasis -88.93 8.27
Sustainability * 8.6 0.95
Food-Drink Homeostasis -175.05 16.07
Food-Drink Homeostasis, Gold -170.3 16.91
Food-Drink Homeostasis, Gold-Silver -168.98 17.47
Food Sharing -42.5 20.75
* The “Handwritten Rules Agent” performs worse than “Random Agent” in the Sustainability environment because the random agent takes some time to find the food for the first time, and that allows the food to grow all over the map. The “Handwritten Rules Agent” has the same code for all environments and this code currently does not contain sufficient sustainability considerations. Instead it follows food consumption and food homeostasis objectives. Note that the “Handwritten Rules Agent” still gets a positive score though.
Figure 11: Baseline results as a bar chart
@@ 4 Conclusion and future plans
The purpose of this work is to add to the discourse on evaluating AI safety in a more rigorous fashion. We have extended on previous work through an ensemble of biologically and economically relevant multi-agent, multi-objective environments and proto-cooperative tests.
@@ 4.1 Future plans
4.1.1 New agents
Our next steps include implementing and testing PPO and DQN to include them in the baselines.
Equally importantly, we plan to develop an interface to LLMs in order to measure LLM agents’ performance on the benchmarks as a baseline for comparing against alternatives.
@@ 4.1.2 New environments
After that we plan to proceed to experimenting with more complex benchmarks, focusing on themes of treacherous turn, cooperation, interruptibility, and side-effects, among other related problems. Currently, at least 12 new experiments are in the planning phase, some are already implemented.
We’re also planning on expanding the complexity of the environment to produce confounding factors. These can be added dynamically at training and test time to test the robustness of the model. If an AGI is published, we can then create an extended sandbox for benchmarking that works in a more realistic environment. We plan to do this via the following noisy variables:
@@ 4.1.3 Some of the themes of planned future environments
Tolerance of changes in agent’s objectives caused by certain other agents. This scenario consists of agents having arbitrary hierarchies, where superiors can change the objectives of the subordinates, while the subordinates could obstruct or trigger these changes. This is inspired from Soares et al. (2015).
@@
Tolerance of changes in environment caused by other agents. One of the related scenarios is a warehouse or factory with restricted space and limited amount of shared tools, while each agent has a separate objective. This is loosely inspired from Orseau and Armstrong (2016). While in the original paper the concept referred to the agent’s ability to be stopped or paused, we are redefining the concept of “interruptions” and “interruptibility” to relate only to changes external to the agent such that the agent’s own objectives and operation mode do not necessarily change. This redefinition is made because the earlier introduced concept of “corrigibility” covers the original meaning of the word “interruptibility” as well. We find that we need to better differentiate between agent-internal and agent-external changes, similarly to how in cybernetics and control theory there are two complementary concepts: set-point or target value, versus actual measured value, both of which can change.
@@
Boundaries and buffer zones between groups of agents with different objectives. Breaking the vase scenario is a classical example - in a multiagent variation, the vase would need to be protected only if some agents use it. Another example is agents causing fires in a forest while working there. They need to stop the fire from spreading to the territory of other agents.
@@
Diminishing returns both in performance and safeguarding objectives of each agent. Turn-taking. One example is a scenario where there are two or more types of shared resources in the environment. Resources can be accessed by one agent at a time. Each agent needs to collect both types of resources.
@@
Scenario and event handler for an imbalanced power dynamic, where the tested agent is monitored whether they will seize a better reward when the opportunity rises, at the expense of another agent.

### v2

@@ 3.1 Baseline agents
As baselines, we provide the following agent types: - "Random Agent" and "Handwritten Rules Agent" - To establish approximate bounds of the lowest and highest expected scores per each benchmark. - DQN, PPO, and A2C - Using OpenAI Stable Baselines 3 algorithm implementations. Each of the aforementioned algorithms was run in three configurations: CNN with 2 layers, CNN with 3 layers, and MLP. - LLM agent - Currently using gpt-4o-mini with OpenAI API.
@@ 3.1.1 Random Agent (no learning)
This agent type represents the approximate lowest scores an incompetent agent is expected to get in the benchmarks (unless it actively takes extreme actions). This agent does not have a capability to learn.
@@ 3.1.2 Handwritten Rules Agent (no learning)
This agent type represents the approximate best scores an agent is expected to achieve on the benchmarks. This agent is based on the simulation of hand-written heuristic action rules for a particular benchmark setup. This agent does not have an actual learning capacity.
3.1.3 OpenAI Stable Baselines 3 algorithms
We used OpenAI implementations, except for the CNN-based feature detector part, where we had to use our custom CNN class, because the OpenAI implementation required a bigger input image than our environment provided.
Because OpenAI Baselines 3 requires a Gym API interface, which does not support multiple agents, for multi-agent scenarios, we created a wrapper that runs each agent in a subprocess with a Gym interface. The wrapper is available in a public repository  [^3] https://github.com/levitation-opensource/zoo_to_gym_multiagent_adapter and is automatically installed as a dependency when setting up the benchmarks Python environment.
For testing PPO on the food sharing benchmark, we used two training modes: either with weight sharing between agents or without weight sharing, in which case the agents had independent models.
3.1.4 LLM agent
Currently using gpt-4o-mini with OpenAI API. Support for other API-s can be easily added in the future.
There is a system prompt describing the general setup. During each step, observations, interception metrics, and rewards are encoded into a text format. Then the model responds with a one-word action choice. The messages are appended to a list so that the model has access to the message history. When the context window becomes full, the excess messages are automatically dropped from the beginning of the list, while the system prompt is preserved.
During training, the observations provided to the model included rewards. During the test, the rewards were not revealed to the model.
@@ 3.2 Experimental results
Since the above-mentioned baseline agents do not learn, the benchmarks did not contain a training phase. We measured the performance only during the testing phase. The size of the environment map was 7 tiles horizontally and vertically. Of these 7x7 maps the outer ring consisted of walls, which means the inner map for actual activities had a size of 5x5 tiles.
The "Random Agent" and "Handwritten Rules Agent" simulations were tested for 100 episodes, each episode consisting of 400 steps.
OpenAI Stable Baselines 3 algorithms were trained for 2500 episodes and tested for 100 episodes, each episode consisting of 400 steps. Due to compute constraints, there were no trials (model resets).
Due to budget reasons, the LLM agent was trained for only 1 episode and then tested also for 1 episode per each benchmark, again with no trials.
In the following table, you can find the average score per test step for each type of agent.
@@
Benchmark Random
DQN
2 CNN
Layers
DQN
3 CNN
Layers
DQN
MLP
PPO
2 CNN
Layers
PPO
3 CNN
Layers
PPO
MLP
A2C
2 CNN
Layers
A2C
3 CNN
Layers
A2C
MLP
GPT-
4o-
mini
Hand-
written
Rules
Food
Unbounded
0.85 19.88 19.88 19.88 11.74 6.38 14.53 1 1 0.6 19.95 19.9
Danger Tiles -1.15 19.88 19.87 19.49 1.78 5.87 -1.11 -1.79 0 -1.6 19.85 19.89
Predators -3.81 19.2 19.19 19.09 -47.72 -47.72 -29.58 -52.61 -50.24 1.87 -67.25 18.19
Homeostasis -90.75 -92.2 -92.04 -90.77 -92.04 -92.2 -92.55 -92.75 -92.04 -92.75 -92.75 8.2
Sustainability 9.04 1.53 2.79 0.84 0.59 1.59 3.45 0.18 1.01 0.14 1.05 0.64 *
Food-Drink
Homeostasis
-181.76 -183.93 -185.5 -181.25 -183.93 -184.6 -184.24 -184.23 -184.23 -184.23 -185.5 16.12
Food-Drink
Homeostasis,
Gold
-177.25 -182.6 -182.76 -167.79 -182.79 -182.37 -182.75 -182.6 -185.5 -185.5 -169.05 16.73
Food-Drink
Homeostasis,
Gold-Silver
-176.47 -183.86 -182.3 -176.7 -182.14 -181.97 -182.16 -183.86 -182.13 -182.13 -167.6 16.82
Food Sharing -43.72 -40.9 -46.38 -43.68
-45.25
-44.42
-45.25
-43.1
-45.25
-46.08
-44.97 -43.64 -44.97 5.67 20.75
* The “Handwritten Rules Agent” performs worse than the "Random Agent" in the sustainability environment because the random agent takes some time to find the food for the first time and that allows the food to grow all over the map. The “Handwritten Rules Agent” has the same code for all environments and currently does not contain sufficient sustainability considerations. Instead, it follows the objectives of food consumption and food homeostasis. However, note that the ’Handwritten Rules Agent’ still gets a positive score.
** PPO with weight sharing
*** PPO without weight sharing
@@ 4 Conclusion and future plans
The purpose of this work is to add to the discourse on evaluating AI alignment and safety in a more rigorous way. We have extended on previous work through an ensemble of biologically and economically relevant multi-agent, multi-objective environments and proto-cooperative tests.
Suffice to say, if throughout a statistically significant number of runs an RL or LLM agent would fail even once, there might be considerable risks in long-term deployment of the said agent in the real world. In our preliminary experiments with multi-trial DQN learning, we saw, for example, cases where the agent performed notably well in 95% of cases, yet failed badly in 5% of cases.
@@ 4.1 Future plans
4.1.1 Benchmarking more LLM models
Our next steps include running the existing benchmarks and our implementation of the LLM agent on more models to obtain the baseline results of different language models.
@@ 4.1.2 New environments
After that, we plan to proceed to experimenting with more complex benchmarks, focusing on themes of treacherous turn, cooperation, interruptibility, and side effects, among other related problems. Currently, at least 12 new experiments are in the planning phase, some have already been implemented.
We’re also planning on expanding the complexity of the environment to produce confounding factors. These can be dynamically added at training and test time to test the robustness of the model. If an AGI is published, we can then create an extended sandbox for benchmarking that works in a more realistic environment. We plan to do this through the following noisy variables:
@@ 4.1.3 Some of the themes of planned future environments
Tolerance to changes in the objectives of the agent caused by certain other agents. This scenario consists of agents having arbitrary hierarchies, where superiors can change the objectives of the subordinates, while the subordinates could obstruct or trigger these changes. This is inspired by Soares et al. (2015).
@@
Tolerance to changes in the environment caused by other agents. One of the related scenarios is a warehouse or factory with restricted space and a limited number of shared tools, while each agent has a separate objective. This is loosely inspired by Orseau and Armstrong (2016). Although in the original paper the concept referred to the agent’s ability to be stopped or paused, we are redefining the concept of “interruptions” and “interruptibility” to relate only to changes external to the agent such that the agent’s own objectives and operation mode do not necessarily change. This redefinition is made because the earlier introduced concept of “corrigibility” also covers the original meaning of the word “interruptibility”. We find that we need to better differentiate between agent-internal and agent-external changes, similarly to how in cybernetics and control theory there are two complementary concepts: set-point or target value, versus actual measured value, both of which can change.
@@
Boundaries and buffer zones between groups of agents with different objectives. Breaking the vase scenario is a classical example - in a multi-agent variation, the vase would need to be protected only if some agents use it. Another example is agents causing fires in a forest while working there. They must stop the fire from spreading to the territory of other agents. This is inspired by Amodei et al. (2016).
@@
Diminishing returns both in performance and safeguarding objectives of each agent. Turn-taking. An example is a scenario where there are two or more types of shared resources in the environment. Resources can be accessed by one agent at a time. Each agent needs to collect both types of resources.
@@
Scenario and event handler for an imbalanced power dynamic, where the tested agent is monitored whether they will seize a better reward when the opportunity rises, at the expense of another agent. This is inspired by Bostrom (2014).
Tiling agents
The environment would use the same homeostatic, diminishing returns, sustainability, and resource sharing objectives as in our existing benchmarks, as well as any new objectives from the new benchmarks described above.
The novelty would be that the main agent would be able to spawn subagents. The subagents would be hard-coded NPC agents. There would be two types of subagents: type A would be aligned, the other type B would be a utility monster. Our main agent can choose whether to produce subagents of type A or type B. It then needs to learn which agent type is aligned and produce only these.
Later we can develop this further for LLM-s so that the sub-agents are not hardcoded. Instead, the LLM agent should reproduce its own system prompt or other training summary and provide it to the sub-agents. If the system prompt or transferable training information gets corrupted during spawning, then the sub-agents may become unaligned.
This theme is inspired by Yudkowsky and Herreshoff (2013).
@@ 5 Acknowledgements
The authors thank Peter van Doesburg for sharing inspiring perspectives and knowledge, which has turned out to be invaluable.